# On the Casimir number and formal codegree of Haagerup-Izumi fusion rings

### v1

@@ 1 Introduction
The discovery of the Haagerup fusion category is one the results in the classification of subfactors of small index [1, 4, 5]. Explicitly, the fusion rule of Haagerup fusion category is as follows:
@@
Surprisingly, currently we still do not know whether the Haagerup fusion category can be obtained from representations of quantum groups at root of unity. Meanwhile, with the help of the Cuntz algebra, M. Izumi generalized the Haagerup fusion rules to a family of fusion rings determined by an arbitrary finite abelian group [5], which is called the Haagerup-Izumi fusion rings now.
Generally, let $G$ be a finite group. A Haagerup-Izumi fusion ring $\mathcal{HI}_{G}$ is a fusion ring of rank (i.e., the cardinal of the $\mathbb{Z}_{+}$-basis of $R$) $2|G|$, with generators $\{g,gX|g\in G\}$ satisfying the following relations
@@
On the other hand, the Casimir operator for Frobenius algebras was introduced in [6], and the author studied the Casimir operator for both $H$ and its Grothendieck ring of $\text{Rep}(H)$, where $H$ is a Hopf algebra and $\text{Rep}(H)$ is the category of finite-dimensional representations of $H$. In [11], the authors defined the Casimir number and determinant of fusion rings, and they investigated relations between these concepts and related concepts, such as the global dimension and Frobenius-Schur exponent of fusion categories, see [3] for specific definition. It was asserted in [11] that the Casimir number and the determinant of the Grothendieck ring of a fusion category share the same set of prime factors. By determining the Casimir number and determinant of Haagerup-Izumi fusion rings $\mathcal{HI}_{\mathbb{Z}_{n}}$, we find that their statement is incorrect, see Proposition 3.1 and Proposition 3.3. Indeed, in the proof of [11, Proposition 2.2], the authors used the commutativity of fusion rules inattentively.
@@ 2 Preliminary
Let $\mathcal{C}$ be a pivotal fusion category with pivotal structure $j$, then categorical dimensions $\dim(-)$ determined by $j$ induces a homomorphism from $\text{Gr}(\mathcal{C})$ to $\mathbb{C}$. If $\dim_{j}(X)=\dim_{j}(X^{*})$ is true for all objects $X$ of $\mathcal{C}$, then $\mathcal{C}$ is spherical [3]. Then the global dimension of $\mathcal{C}$ is defined as
@@
The element $c(\mathbf{1})=\sum_{X\in\mathcal{O}(\mathcal{C})}XX^{*}$ is called the Casimir element of $\text{Gr}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $\mathcal{C}$. The intersection $\mathbb{Z}\cap\text{Im}(c)$ is a nonzero principal ideal of $\mathbb{Z}$, and the positive generator of $\mathbb{Z}\cap\text{Im}(c)$ is called the Casimir number of $\mathcal{C}$. Observe that the matrix $c(\mathbf{1})]$ of left multiplication by $c(\mathbf{1})$ with respect to the basis of $\text{Gr}(\mathcal{C})$ is a positive definite integer matrix [6, Proposition 8], and its determinant $\text{det}[c(\mathbf{1})]$ is called the determinant of $\mathcal{C}$.
@@ 3.1 The Casimir number and determinant
It follows from the definition of Casimir number that
@@
Therefore, $\lambda_{n}+...+\lambda_{2n-1}=-\frac{2n}{n^{2}+n+4}\lambda_{0}$, it follows that $c(x)=\frac{2n(n^{2}+4)}{n^{2}+n+4}\lambda_{0}\mathbf{1}$. Since $n$ is even, $(n^{2}+n+4,n)=\left\{\begin{array}[]{ll}4,&\hbox{if $4\mid n$}\\ 2,&\hbox{if $4\nmid n$.}\end{array}\right$.. Same as Proposition 3.1, the Casimir number of Haagerup-Izumi fusion ring $\mathcal{HI}_{\mathbb{Z}_{n}}$ is $\frac{n(n^{2}+4)}{2}$ if $4\mid n$, otherwise, it is $(n^{2}+4)n$.∎
@@
If $n$ is odd, Theorem 3.1 states that the Casimir number of $\mathcal{HI}_{\mathbb{Z}_{n}}$ is $n(n^{2}+4)$, which is odd too, so it does not share the same set of prime divisors with the determinant of $\mathcal{HI}_{\mathbb{Z}_{n}}$ by Proposition 3.3. Note that there exist fusion categories with Haagerup-Izumi fusion rules for odd integers less than $31$ [2], therefore, this provides a counterexample to [11, Corollary 2.4]. Indeed, in the proof of [11, Proposition 2.2], the authors used the commutativity of the fusion rules of fusion categories implicitly.
@@ 3.2 Irreducible representations and formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$
If $n=1,2$, it is easy to see that homomorphisms in Lemma 3.5 are all the isomorphism classes of irreducible representations of $\mathcal{HI}_{\mathbb{Z}_{n}}$. We assume $n\geq 3$ below.
@@
Let $M_{i}:=\left({\begin{array}[]{cc}\zeta_{n}^{i}&0\\ 0&\zeta_{n}^{-i}\end{array}}\right)$ and $N_{i}:=\left({\begin{array}[]{cc}0&1\\ 1&0\end{array}}\right)$. A direct computation shows that these matrices satisfies the generating Equations 2, so the $A$-module structure is well-defined. ∎
@@
Let $V$ be an irreducible representation of $A$ with $\dim_{\mathbb{C}}(V)>1$. Then $V$ is isomorphic to one of the $V_{i}$ defined by Equations (5), $1\leq i\leq\big{[}\frac{n-1}{2}\big{]}$.
@@
Let $n\geq 3$ be an odd integer, the other case is similar. A direct computation shows $c(\mathbf{1})=2n\mathbf{1}+n\sum_{j=n}^{2n-1}Y_{j}$. Then Lemma 3.5 shows that $c(\mathbf{1})$ acts on one-dimensional representations as scalar $\text{FPdim}(\mathcal{HI}_{\mathbb{Z}_{n}})$ and $\sigma(\text{FPdim}(\mathcal{HI}_{\mathbb{Z}_{n}}))$, which are formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$. And $c(\mathbf{1})$ acts on all two-dimensional irreducible representations as $2n$ by Proposition 3.7, so all the corresponding formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$ are $n$ by definition. Meanwhile,
@@ Acknowledgements
The authors thanks L. Li for discussions. The third author was supported by NSFC (no. 12571041) and Qinglan Project of Yangzhou University.

### v2

@@ 1 Introduction
The discovery of the Haagerup fusion category is one of the results in the classification of subfactors of small index [1, 4, 5]. Explicitly, the rank (i.e., the number of isomorphism classes of simple objects) of the Haagerup fusion category is $6$, and its fusion rule is as follows:
@@
Surprisingly, currently we still do not know whether the Drinfeld center of the Haagerup fusion category can be obtained from the semisimplification of tensor categories of representations of quantum groups at root of unity. Meanwhile, with the help of the Cuntz algebra, M. Izumi generalized the Haagerup fusion rules to a family of fusion rings determined by an arbitrary finite abelian group [5], which is called the Haagerup-Izumi fusion rings now.
Generally, let $G$ be a finite group. The Haagerup-Izumi fusion ring $\mathcal{HI}_{G}$ is a fusion ring with $\mathbb{Z}_{+}$-basis $\{g,gX|g\in G\}$ and they satisfy the following relations
@@
On the other hand, the Casimir operator for Frobenius algebras was introduced in [6], and the author studied the Casimir operator for both $H$ and the Grothendieck ring of $\text{Rep}(H)$, where $H$ is a Hopf algebra and $\text{Rep}(H)$ is the category of finite-dimensional representations of $H$. In [12], the authors defined the Casimir number and determinant of fusion rings, and they investigated relations between these concepts and related concepts, such as the global dimension and Frobenius-Schur exponent of fusion categories, see [3] for specific definition. It was asserted in [12] that the Casimir number and the determinant of the Grothendieck ring of a fusion category share the same set of prime factors. By determining the Casimir number and determinant of Haagerup-Izumi fusion rings $\mathcal{HI}_{\mathbb{Z}_{n}}$, we find that their statement is incorrect, see Proposition 3.1 and Proposition 3.3. Indeed, in the proof of [12, Proposition 2.2], the authors used the commutativity of fusion rules inattentively.
@@ 2 Preliminary
Let $\mathcal{C}$ be a pivotal fusion category with a pivotal structure $j$, then the categorical dimensions $\dim(-)$ determined by $j$ induces a homomorphism from $\text{Gr}(\mathcal{C})$ to $\mathbb{C}$. If $\dim_{j}(X)=\dim_{j}(X^{*})$ is true for all objects $X$ of $\mathcal{C}$, then $\mathcal{C}$ said to be spherical [3]. And the global dimension of $\mathcal{C}$ is defined as
@@
The element $c(\mathbf{1})=\sum_{X\in\mathcal{O}(\mathcal{C})}XX^{*}$ is called the Casimir element of $\text{Gr}(\mathcal{C})$, where $\mathbf{1}$ is the unit object of $\mathcal{C}$. The intersection $\mathbb{Z}\cap\text{Im}(c)$ is a nonzero principal ideal of $\mathbb{Z}$, and the positive generator of $\mathbb{Z}\cap\text{Im}(c)$ is called the Casimir number of $\mathcal{C}$ [12]. It is well-known that the matrix $c(\mathbf{1})]$ of left multiplication by $c(\mathbf{1})$ with respect to the basis of $\text{Gr}(\mathcal{C})$ is a positive definite integer matrix, and its determinant $\text{det}[c(\mathbf{1})]$ is called the determinant of $\mathcal{C}$ [12].
@@ 3.1 The Casimir number and determinant
It follows from the definition of the Casimir number that
@@
Therefore, $\lambda_{n}+...+\lambda_{2n-1}=-\frac{2n}{n^{2}+n+4}\lambda_{0}$, it follows that $c(x)=\frac{2n(n^{2}+4)}{n^{2}+n+4}\lambda_{0}\mathbf{1}$. Since $n$ is even, the largest common divisor $(n^{2}+n+4,n)=\left\{\begin{array}[]{ll}4,&\hbox{if $4\mid n$}\\ 2,&\hbox{if $4\nmid n$.}\end{array}\right$.. By using a similar argument as Proposition 3.1, the Casimir number of Haagerup-Izumi fusion ring $\mathcal{HI}_{\mathbb{Z}_{n}}$ is $\frac{n(n^{2}+4)}{2}$ if $4\mid n$, otherwise, it is $(n^{2}+4)n$.∎
@@
If $n$ is odd, Theorem 3.1 states that the Casimir number of $\mathcal{HI}_{\mathbb{Z}_{n}}$ is $n(n^{2}+4)$, which is odd too, so it does not share the same set of prime divisors with the determinant of $\mathcal{HI}_{\mathbb{Z}_{n}}$ by Proposition 3.3. Note that there exist fusion categories with Haagerup-Izumi fusion rules for odd integers less than $31$ [2], therefore, this provides a counterexample to [12, Corollary 2.4]. Indeed, in the proof of [12, Proposition 2.2], the authors used the commutativity of the fusion rules of fusion categories implicitly.
@@ 3.2 Irreducible representations and formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$
If $n=1,2$, then $\mathcal{HI}_{\mathbb{Z}_{n}}$ is commutative, it is easy to see that homomorphisms in Lemma 3.5 are all the isomorphism classes of irreducible representations of $A$. We assume $n\geq 3$ below.
@@
Let $M_{i}:=\left({\begin{array}[]{cc}\zeta_{n}^{i}&0\\ 0&\zeta_{n}^{-i}\end{array}}\right)$ and $N_{i}:=\left({\begin{array}[]{cc}0&1\\ 1&0\end{array}}\right)$, where $1\leq i\leq\big[\frac{n-1}{2}\big]$. A direct computation shows that these two matrices satisfies the generating the Equations 2, so the $A$-module structure is well-defined. ∎
@@
Let $V$ be an irreducible representation of $A$ with $\dim_{\mathbb{C}}(V)>1$. Then $V$ is isomorphic to one of the $V_{i}$ defined by the Equations (5), $1\leq i\leq\big[\frac{n-1}{2}\big]$.
@@
We note that irreducible representations for fusion rings with two orbits are also constructed in [11, Proposition 2.16].
@@
Let $n\geq 3$ be an odd integer, the other case is similar. A direct computation shows
$c(\mathbf{1})=2n\mathbf{1}+n\sum_{j=n}^{2n-1}Y_{j}.$
Then Lemma 3.5 shows that $c(\mathbf{1})$ acts on one-dimensional representations as scalar $\text{FPdim}(\mathcal{HI}_{\mathbb{Z}_{n}})$ and $\sigma(\text{FPdim}(\mathcal{HI}_{\mathbb{Z}_{n}}))$, which are formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$. And $c(\mathbf{1})$ acts on all two-dimensional irreducible representations as $2n$ by Proposition 3.7, so all the corresponding formal codegrees of $\mathcal{HI}_{\mathbb{Z}_{n}}$ are $n$ by definition. Meanwhile,
@@ Acknowledgements
The authors thanks L. Li and A. Schopieray for comments. The third author was supported by NSFC (no. 12571041) and Qinglan Project of Yangzhou University.